OpenStudy (anonymous):

Please I really need help(Logarithms): The weight of bacteria in a culture t hours after it has been established is given by W=2.5x2^0.04t grams. After what time will the weight reach 4 grams?

OpenStudy (anonymous):

I tried solving it like this: log4=log2.5 x 0.04t x log2 And then I just solved the equation but I keep getting it wrong

OpenStudy (anonymous):

Thanks for your help :) I'll just check if it's right

OpenStudy (anonymous):

okay the thing is we are just supposed to use base 10

OpenStudy (anonymous):

and the answer is 17 but I get something else

OpenStudy (anonymous):

\[2.5\times 2^{0.04t}=5\] \[2^{0.04t}=2\] \[.04t=1\] \[t=1\div.04\]

OpenStudy (anonymous):

oooh it said 4, not 5, my mistake

OpenStudy (anonymous):

\[2.5\times 2^{.04t}=4\] \[2^{.04t}=1.6\] \[.04t=\frac{\ln(1.6)}{\ln(2)}\] \[t=\frac{\ln(1.6)}{.04\ln(2)}\]

OpenStudy (anonymous):

If you're interested in knowing the correct solution:

OpenStudy (anonymous):

4=2.5x2^0.04t Using the rules of the logarithms to expand this equation gives after a few steps log4 = log2.5 + 0.04t x log2 Now solve this for t to get t=(log4 - log2.5)/(0.04xlog2) = 16.95 or approx. 17.0 h